OpenStudy (anonymous):

is radical 7 times radical 7 radical 49

OpenStudy (jamesj):

Yes

OpenStudy (anonymous):

thanks, so is radical 3 times radical 3 radical 9

OpenStudy (jamesj):

Yes. By definition "radical 3" means the square root of 3. In symbols\[\sqrt{3} = 3^{1/2}\] Hence \[\sqrt{3} \sqrt{3} = 3^{1/2} 3^{1/2} = (3 \times 3)^{1/2} = 9^{1/2} = \sqrt{9}\]

OpenStudy (jamesj):

which makes complete sense because \[\sqrt{9} = 3\] and it had better be the case that \[\sqrt{3} \sqrt{3} = \sqrt{3^2} = 3\]